PLEASE HELP ME !!!NEED STEP BY STEP SOLUTION !!THANK'S, JERRY GARRETT
2+4/a=14/a-3 ????

Answer by stanbon(75887)

2+4/a=14/a-3
Without parenthese it is hard to tell where numerators and denominators
start and end.
Here's a guess on that:
2 + (4/a) = (14/a)-3
Subtract 2 from both sides to get:
(4/a) = (14/a)-5
Subtract (14/a) from both sides to get:
-10/a = -5
Invert both sides to get:
a/-10 = -1/5
Multiply both sides by -10 to get:
a= 2
------------
Cheers,
Stan H.